# The Limited Impact of Medical Adaptation of Large Language and Vision-Language Models

### v1

@@ Abstract
Several recent works seek to develop foundation models specifically for medical applications, adapting general-purpose large language models (LLMs) and vision-language models (VLMs) via continued pretraining on publicly available biomedical corpora. These works typically claim that such domain-adaptive pretraining (DAPT) improves performance on downstream medical tasks, such as answering medical licensing exam questions. In this paper, we compare ten public “medical” LLMs and two VLMs against their corresponding base models, arriving at a different conclusion: all medical VLMs and nearly all medical LLMs fail to consistently improve over their base models in the zero-/few-shot prompting and supervised fine-tuning regimes for medical question-answering (QA). [^1] This working paper substantially extends the results of Jeong et al. (2024) to include closed-ended tasks based on clinical notes in addition to medical-exam-style QA, as well as a comparison of performance when using medical versus general-domain models as an initialization for supervised fine-tuning. For instance, across all tasks and model pairs we consider in the 3-shot setting, medical LLMs only outperform their base models in 22.7% of cases, reach a (statistical) tie in 36.8% of cases, and are significantly worse than their base models in the remaining 40.5% of cases. Our conclusions are based on (i) comparing each medical model head-to-head, directly against the corresponding base model; (ii) optimizing the prompts for each model separately in zero-/few-shot prompting; and (iii) accounting for statistical uncertainty in comparisons. While these basic practices are not consistently adopted in the literature, our ablations show that they substantially impact conclusions. Meanwhile, we find that after fine-tuning on specific QA tasks, medical LLMs can show performance improvements, but the benefits do not carry over to tasks based on clinical notes. Our findings suggest that state-of-the-art general-domain models may already exhibit strong medical knowledge and reasoning capabilities, and offer recommendations to strengthen the conclusions of future studies.
Keywords: domain-adaptive pretraining, large language models (LLMs), vision-language models (VLMs), medical question-answering
@@ 1 Introduction
To address such concerns, recent works have proposed cheaper, open-source alternatives through domain-adaptive pretraining (DAPT; Gururangan et al., 2020), where a pretrained open-source general-domain model—such as Llama (Touvron et al., 2023a, b; Meta, 2024) or Mistral (Jiang et al., 2023) in the language space, and LLaVA (Liu et al., 2023) or Open-Flamingo (Awadalla et al., 2023) in the vision-language space—is continually pretrained on biomedical (image-)text corpora from public sources such as PubMed and medical textbooks. While some prior works show that medical models pretrained from scratch only using domain-specific corpora can outperform those trained via DAPT, both in the context of BERT-style encoder-only models (Devlin et al., 2019; Gu et al., 2021; Yang et al., 2022) and decoder models (Taylor et al., 2022; Luo et al., 2022; Hernandez et al., 2023; Bolton et al., 2024), the DAPT approach has become common practice, resulting in a trend where the release of a more capable general-domain model is typically followed by the release of its medical counterpart.
@@
Despite the widespread adoption of medical DAPT, the claimed improvements in performance are worth scrutinizing. While the story is intuitive, more recent base models often already exhibit strong off-the-shelf performance on medical benchmarks without any adaptation. For instance, as of the time of writing, the general-domain Llama-3-8B (Meta, 2024) outperforms other medically specialized models such as MediTron-70B (Chen et al., 2023) and BioMistral-7B (Labrak et al., 2024) on the Open Medical LLM Leaderboard (Pal et al., 2024), which evaluates each model on standard medical QA benchmark datasets such as MedQA (Jin et al., 2020) and MedMCQA (Pal et al., 2022). Moreover, given the general lack of transparency about the pretraining corpora used to train the general-domain model in the first place, it is possible that they may already be trained on relevant medical text.
@@
Perhaps more concerning is the lack of apples-to-apples comparisons in the literature. First, medical models resulting from DAPT are often only compared against other baselines with different architectures and model scale (e.g., Clinical-Camel-70B (Toma et al., 2023) vs. GPT-4 (OpenAI, 2023a)). Second, even for the same model scale, models are often evaluated under inconsistent evaluation setups (e.g., MediTron-70B (Chen et al., 2023) fine-tuned on MedQA with gradient updates vs. Clinical-Camel-70B (Toma et al., 2023) few-shot prompted on MedQA). Third, the common practice of using a single, fixed prompting setup (e.g., prompt format, choice of few-shot examples) for all models under evaluation also warrants concern, as LLM/VLM behavior is extremely sensitive to such design decisions (Jiang et al., 2020; Zhao et al., 2021; Ceballos-Arroyo et al., 2024b), and the “optimal” choice of such details rarely correlates between different models (Sclar et al., 2024). All of these issues can confound the interpretation of results, when evaluating the performance benefits from medical DAPT.
Figure 1: Medical LLMs and VLMs trained via domain-adaptive pretraining (DAPT) show limited improvement over their general-domain counterparts in zero-/few-shot prompting and supervised fine-tuning regimes. (a) Overview of our head-to-head evaluation approach for each pair of general-domain (blue) and medically adapted LLM/VLM (red). (b) Win, tie, and loss rates (%) of medical models vs. their corresponding base models across all (model pair, QA dataset) combinations. Win rate refers to the proportion of (model pair, QA dataset) combinations where a medical model shows a statistically significant improvement.
@@
In this paper, we perform an apples-to-apples comparison that addresses these concerns, comparing ten medical LLMs and two medical VLMs against their general-domain base models on various medical (visual) QA tasks (see Figure 1(a)). We compare several pairs of general-domain and medically adapted LLMs/VLMs (see Table 1), whose only differences lie in medical DAPT (i.e., one model is the base model, from which the other is derived via medical DAPT). For each pair, we compare their downstream performances from (i) zero-/few-shot prompting (Radford et al., 2019; Brown et al., 2020) and (ii) supervised fine-tuning. For the former, we follow the approach considered in Jeong et al. (2024) and compare the performances after independently selecting the “best” prompt format and few-shot examples for each model based on the validation set (Section 3.1). For the latter, we compare the performances after fine-tuning each model on the training set of each downstream QA dataset, using the best hyperparameters selected via grid search (Section 3.2). In both cases, we use the percentile bootstrap to assess whether the perceived improvements in performance from medical DAPT are attributable to chance.
LLM Llama-3-70B-Instruct (Meta, 2024) Med42-v2-70B (Christophe et al., 2024b) Medical QA Datasets (e.g., MedQA, MedMCQA)
Medical Instruction 120k (Altaf, 2023)
OpenGPT (OpenChat) (Wang et al., 2024)
StackExchange (Lambert et al., 2023)
Medical Flashcards (Han et al., 2023)
Llama-3-70B-Instruct (Meta, 2024) OpenBioLLM-70B (Pal and Sankarasubbu, 2024) Undisclosed
Llama-2-70B (Touvron et al., 2023b) MediTron-70B (Chen et al., 2023) Clinical Practice Guidelines (e.g., CDC, WHO)
PubMed Articles (S2ORC; Lo et al., 2020)
Llama-2-70B (Touvron et al., 2023b) Clinical-Camel-70B (Toma et al., 2023) ShareGPT
20k PubMed Articles Published Before 2021
Random 4k Subset of MedQA (Jin et al., 2020)
Llama-2-70B (Touvron et al., 2023b) Med42-v1-70B (Christophe et al., 2024a) Medical QA Datasets (e.g., MedQA, MedMCQA)
OpenGPT (OpenChat) (Wang et al., 2024)
StackExchange (Lambert et al., 2023)
Medical Flashcards (Han et al., 2023)
CORD-19 (Wang et al., 2020)
Llama-3-8B-Instruct (Meta, 2024) Med42-v2-8B (Christophe et al., 2024b) Medical QA Datasets (e.g., MedQA, MedMCQA)
Medical Instruction 120k (Altaf, 2023)
OpenGPT (OpenChat) (Wang et al., 2024)
StackExchange (Lambert et al., 2023)
Medical Flashcards (Han et al., 2023)
Llama-3-8B (Meta, 2024) OpenBioLLM-8B (Pal and Sankarasubbu, 2024) Undisclosed
Llama-2-7B (Touvron et al., 2023b) MediTron-7B (Chen et al., 2023) Clinical Practice Guidelines (e.g., CDC, WHO)
PubMed Articles (S2ORC; Lo et al., 2020)
Mistral-7B-Instruct-v0.1 (Jiang et al., 2023) BioMistral-7B (Labrak et al., 2024) PubMed Articles (PMC Open Access Subset)
Llama-2-7B-Chat (Touvron et al., 2023b) BioMedGPT-LM-7B (Luo et al., 2023) PubMed Articles (S2ORC; Lo et al., 2020)
VLM LLaVA-v0-7B (Liu et al., 2023) LLaVA-Med-7B (Li et al., 2023a) PubMed Articles (PMC-15M; Zhang et al., 2023)
Open-Flamingo-9B (Awadalla et al., 2023) Med-Flamingo-9B (Moor et al., 2023b) Medical Textbooks (MTB; Moor et al., 2023b)
PubMed Articles (PMC-OA; Lin et al., 2023)
@@
Overall, we find that both medical LLMs and VLMs show limited improvement over their base models, across all medical QA tasks and evaluation settings that we consider (Figure 1(b)). In the zero-/few-shot prompting regime (Section 4), we find that all medical VLMs and the majority of medical LLMs fail to consistently outperform their base models across all datasets, including QA tasks focused on assessing medical knowledge and those based on real-world clinical notes (e.g., discharge summaries). In the supervised fine-tuning regime (Section 5), we find that the medical LLMs overall do show statistically significant improvements on medical knowledge QA tasks but not on the clinical note QA tasks, while the medical VLMs show little to no improvement on all of the visual medical QA tasks. Our findings also suggest that rigorous pairwise comparison of models, including tests for statistical significance, is essential to drawing reliable conclusions about the performance benefits from medical DAPT.
@@
    We provide a comprehensive head-to-head comparison between state-of-the-art general-domain LLMs/VLMs and their medical DAPT counterparts on various medical QA benchmarks, to investigate the effectiveness of DAPT for medical specialization.
@@
    We find that in the supervised fine-tuning regime, all medical VLMs fail to show improvement, while medical LLMs show improvement on textual medical knowledge QA tasks but not on tasks based on clinical notes (Section 5).
@@ 2 Related Work
More recent works suggest that decoder-based autoregressive LLMs and VLMs trained via medical DAPT also show strong performance on various downstream medical tasks. Medical LLMs such as MediTron (Chen et al., 2023), adapted from Llama-2 (Touvron et al., 2023b); and BioMistral (Labrak et al., 2024), adapted from Mistral-7B-Instruct-v0.1 (Jiang et al., 2023); perform well on knowledge-intensive QA tasks derived from medical licensing and academic exams (Jin et al., 2020; Pal et al., 2022; Hendrycks et al., 2021) and PubMed abstracts (Jin et al., 2019). Medical VLMs such as LLaVA-Med (Li et al., 2023a), adapted from LLaVA (Liu et al., 2023); and Med-Flamingo (Moor et al., 2023b), adapted from Open-Flamingo (Awadalla et al., 2023); also perform well on visual QA tasks based on radiology (Lau et al., 2018; Liu et al., 2021) and pathology images (He et al., 2020) and academic exams (Yue et al., 2024). These encouraging results have established DAPT as a go-to approach for training medically specialized LLMs and VLMs, a conclusion that we re-examine in the zero-/few-shot prompting and supervised fine-tuning regimes.
@@ 3 Experimental Setup
To investigate the effectiveness of medical DAPT in improving (i) zero-/few-shot and (ii) supervised fine-tuning performances, we compare ten medical LLMs and two medical VLMs against their general-domain counterparts in pairs (Figure 1(a)), on 22 textual QA datasets (8 based on clinical notes) and 8 visual QA datasets, respectively. The models in each pair are exactly identical in model architecture and scale, and their only difference lies in whether they were additionally pretrained on medical data. We also note that while some of datasets considered contain both closed-ended (i.e., has clear ground-truth answers) and open-ended questions, we focus our evaluations on the former, where an objective, quantitative assessment of medical knowledge and reasoning capabilities is possible. For reproducibility of our results, we open-source the source code used for all of our evaluations described below via our GitHub repository [^2] https://github.com/taekb/eval-medical-dapt.
Models.
In Table 1, we provide a summary of all of the LLM and VLM pairs that we use for evaluation, along with details about the pretraining corpora used for adaptation to the medical domain. For LLaVA (Liu et al., 2023), we use the very first version (v0) that uses Vicuna-v0 (Chiang et al., 2023) as the LLM backbone, as LLaVA-Med (Li et al., 2023a) was adapted from that particular version. For all models, we use the checkpoints made available via HuggingFace. In all zero-/few-shot prompting experiments, we generate predictions from each model via (i) greedy decoding (i.e., sampling with temperature $T=0$) and (ii) constrained decoding. For constrained decoding, we constrain the token vocabulary to be one of the answer choice letters (e.g., one of [“A”, “B”, “C”, “D”] for a four-choice QA dataset) and treat the answer choice with the highest token probability as a given model’s prediction. For the supervised fine-tuning experiments, we only generate predictions via greedy decoding, as it best reflects the setup used for fine-tuning (see Section 3.2).
@@
For textual medical knowledge QA, we use MedQA (Jin et al., 2020), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019), and MMLU-Medical (Hendrycks et al., 2021) for evaluation. MMLU-Medical refers to a subset of MMLU corresponding to 9 subjects related to medicine: anatomy, clinical knowledge, college biology, college medicine, high school biology, medical genetics, nutrition, professional medicine, and virology. For MedQA, we use the official train–validation-test splits as provided through BigBio (Fries et al., 2022b). We note that MedQA has two versions, one with four answer choices per question and the other with five, and we use both for evaluation. For MedMCQA, which does not have a public test set, we follow the approach taken by Wu et al. (2024) and Labrak et al. (2024), taking a random 80–20 train–validation split of the official training set and using the official validation set for testing. For PubMedQA, we follow Singhal et al. (2023a), using the 211k artifically generated QA samples for training, and taking a 50–50 split on the 1k expert-labeled examples. For MMLU-Medical, we use the official split as provided. Meanwhile, we note that as the Med42 models—Med42-v2-70B, Med42-v2-8B, and Med42-v1-70B—were trained on all of the QA datasets mentioned above as part of DAPT, we exclude these models from our discussion on textual medical knowledge QA. We provide the remaining details in Appendix A.
@@
As strong performance on medical knowledge QA tasks may not necessarily imply high clinical utility (Wornow et al., 2023), we also perform evaluations using the following textual QA datasets based on de-identified clinical notes: MedNLI (Romanov and Shivade, 2018), EHRNoteQA (Kweon et al., 2024), the 2008 i2b2 Obesity Comorbidity Detection Challenge dataset (Uzuner, 2009) (4 classification tasks), the CASI Clinical Acronym Sense Disambiguation dataset (Moon et al., 2014), and the MIMIC-III Clinical Acronym Sense Disambiguation dataset (Johnson et al., 2016; Adams et al., 2020). For MedNLI, we frame each natural language inference task as a three-way closed-ended QA task, where the answer to each question is one of [“entailment”, “contradiction”, and “neutral”], and use the official train–validation–test splits provided via PhysioNet (Goldberger et al., 2000) under a data use agreement. For EHRNoteQA, we use the “Level 1” subset, which only contains QA pairs with clinical notes at most 3k tokens in length, to account for the limited context window sizes of the LLMs that we evaluate [^3] Most of the LLMs that we evaluate have a context window size of $\sim$4k tokens. A notable exception is MediTron-7B, which has a context window size of only $\sim$2k tokens. We therefore exclude MediTron-7B and its base model (Llama-2-7B) from the evaluations on EHRNoteQA and the i2b2 datasets., and take a random 60–20–20 split of the data to obtain the train, validation, and test sets. For the 2008 i2b2 Obesity Comorbidity Detection Challenge dataset, we follow the preprocessing steps in Ceballos-Arroyo et al. (2024a) to obtain 4 binary classification (i.e., two-way multiple-choice QA) datasets, each focused on detecting the presence of (i) asthma, (ii) coronary artery disease (CAD), (iii) diabetes, and (iv) obesity from a given clinical note. We then select the QA pairs with clinical notes at most 3k tokens in length, as done for EHRNoteQA. As the preprocessing pipeline by Ceballos-Arroyo et al. (2024a) outputs a train–test split, we take a random 80–20 split on the training set to obtain the train and validation sets. For the CASI and MIMIC-III sense disambiguation datasets, we follow the preprocessing steps detailed in Adams et al. (2020) for consistency with prior works (e.g., Agrawal et al. (2022)). We provide the remaining details in Appendix A.
Visual QA Datasets.

### v2

@@ Abstract
Several recent works seek to adapt general-purpose large language models (LLMs) and vision-language models (VLMs) for medical applications through continued pretraining on publicly available biomedical corpora. These works typically claim that such domain-adaptive pretraining improves performance on various downstream medical tasks, such as answering medical exam questions. In this paper, we compare ten “medical” LLMs and two VLMs against their corresponding base models, arriving at a different conclusion: all medical VLMs and nearly all medical LLMs fail to consistently improve over their base models in the zero-/few-shot prompting and supervised fine-tuning regimes for medical question answering (QA). For instance, on clinical-note-based QA tasks in the 3-shot setting, medical LLMs outperform their base models in only 26.7% of cases, reach a (statistical) tie in 16.7% of cases, and perform significantly worse in the remaining 56.7% of cases. Our conclusions are based on (i) comparing each medical model directly against its base model; (ii) optimizing the prompts for each model separately in zero-/few-shot prompting; and (iii) accounting for statistical uncertainty in comparisons. Our findings suggest that state-of-the-art general-domain models may already exhibit strong medical knowledge and reasoning capabilities, and offer recommendations to strengthen the conclusions of future studies.
Keywords: domain-adaptive pretraining, large language models (LLMs), vision-language models (VLMs), foundation models for medical domain, question answering
@@ 1 Introduction
To address such concerns, recent works have proposed cheaper, open-source alternatives through domain-adaptive pretraining (DAPT; Gururangan et al., 2020), where a pretrained open-source general-domain model—such as Llama (Touvron et al., 2023a, b; Meta, 2024) or Mistral (Jiang et al., 2023) in the language space; and LLaVA (Liu et al., 2023) or Open-Flamingo (Awadalla et al., 2023) in the vision-language space—is continually pretrained on biomedical (image-)text corpora from public sources such as PubMed. While some works show that medical models pretrained from scratch only using domain-specific corpora can outperform those trained via DAPT, both in the context of BERT-style encoder-only models (Devlin et al., 2019; Gu et al., 2021; Yang et al., 2022) and decoder models (Taylor et al., 2022; Luo et al., 2022; Hernandez et al., 2023; Bolton et al., 2024), the DAPT approach has become common practice, resulting in a trend where the release of a more capable general-domain model is typically followed by the release of its medical counterpart.
@@
Despite the widespread adoption of medical DAPT, the claimed improvements in performance are worth scrutinizing. While the central story (improved domain-specific performance through domain-specific adaptation) is intuitive, more recent base models often already exhibit strong off-the-shelf performance on medical benchmarks without any adaptation. For instance, as of the time of writing, the general-domain Llama-3-8B (Meta, 2024) outperforms other medically specialized models such as MediTron-70B (Chen et al., 2023) and BioMistral-7B (Labrak et al., 2024) on the Open Medical LLM Leaderboard (Pal et al., 2024), which evaluates each model on standard medical QA benchmark datasets such as MedQA (Jin et al., 2020) and MedMCQA (Pal et al., 2022). Moreover, given the general lack of transparency about the pretraining corpora used to train the general-domain model in the first place, it is possible that they may already be trained on relevant medical text (or even the very same text as used for medical DAPT).
@@
Perhaps more concerning is the lack of apples-to-apples comparisons in the literature. First, medical models resulting from DAPT are often only compared against other baselines with different architectures and model scale (e.g., Clinical-Camel-70B (Toma et al., 2023) vs. GPT-4 (OpenAI, 2023a)). Second, even for the same model scale, models are often evaluated under inconsistent evaluation setups. For example, Chen et al. (2023) fine-tune MediTron-70B on MedMCQA with gradient updates and compare it to Clinical-Camel-70B (Toma et al., 2023) zero-shot prompted on MedMCQA. Third, the common practice of using a single, fixed prompting setup (e.g., prompt format, choice of few-shot examples) for all models under evaluation also warrants concern, as LLM/VLM behavior is extremely sensitive to such design decisions (Jiang et al., 2020; Zhao et al., 2021; Ceballos-Arroyo et al., 2024), and prompt designs optimized for a proposed method are seldom optimal for the baselines (Sclar et al., 2024). These issues can undermine the validity of claimed performance benefits in the medical DAPT literature.
Meanwhile, medical LLMs trained via DAPT are often solely evaluated on QA tasks designed to assess medical knowledge (e.g., MedQA (Jin et al., 2020), MedMCQA (Pal et al., 2022)), although it is generally unclear whether strong performance on such tasks necessarily implies high clinical utility (Wornow et al., 2023). Notably, clinical notes (e.g., outpatient SOAP notes and discharge summaries) look different from academic biomedical articles (PubMed is the primary source of data used in many medical adaptation papers (Table 2)) both in form, content, and surface-level use of language (e.g., syntactical differences, use of jargon, abundance of naturally occurring grammatical errors (Agrawal et al., 2022; Hernandez et al., 2023)). To assess whether medical DAPT contributes to higher clinical utility, it is essential to also evaluate medical models resulting from DAPT on QA tasks grounded on real-world clinical notes in direct comparison against their base models.
Figure 1: Overview of our head-to-head evaluation approach for each pair of medical LLM/VLM (red) and its base model (turquoise), across a wide variety of medical QA tasks and model use cases. (Left) We compare each model pair based on their zero-/few-shot QA capabilities (top) and their QA performances after supervised fine-tuning (bottom), while ensuring that their only differences lie in medical DAPT. (Right) We substantially extend the results of Jeong et al. (2024) (yellow) to include QA tasks based on clinical notes, as well as a comparison of downstream QA performance when using medical vs. general-domain models as an initialization for fine-tuning (cyan). Within each evaluation regime, a indicates that a model is evaluated in that regime both in the original paper and in our work, while a indicates that it is only evaluated in our work [^2] Although Christophe et al. (2024a, b) describe their evaluation of the proposed Med42 models on medical knowledge QA tasks as being in the “zero-shot” setting, we treat the zero-/few-shot prompting evaluations as missing, since they were already fine-tuned on these datasets via medical DAPT.. None of the medical LLMs we evaluate have previously been evaluated on QA tasks based on real-world clinical notes.
@@
In this paper, we perform an apples-to-apples comparison that addresses these concerns, comparing ten medical LLMs and two medical VLMs against their general-domain base models on various medical (visual) QA tasks (Figure 2). For each pair of general-domain and medically adapted LLMs/VLMs, whose only differences lie in medical DAPT (i.e., one model is the base model, from which the other is derived via medical DAPT), we compare their downstream performances from (i) zero-/few-shot prompting (Radford et al., 2019; Brown et al., 2020) and (ii) supervised fine-tuning (SFT). For the former, we follow Jeong et al. (2024) and compare the performances after independently selecting the “best” prompt format and few-shot examples for each model based on the validation set (Section 3.1). For the latter, we compare the performances after fine-tuning each model on the training set of each downstream QA dataset, using the best hyperparameters selected via grid search (Section 3.2). In both cases, we use the percentile bootstrap to assess whether the perceived improvements in performance from medical DAPT are attributable to chance. In Table 2, we list all of the LLM/VLM pairs included in our study, which were selected to cover a wide range of general-domain base models and model scales (7B–70B).
Figure 2: Medical LLMs and VLMs trained via DAPT show limited improvement over their general-domain counterparts in the zero-/few-shot prompting and supervised fine-tuning (SFT) regimes. Here, we show the win, tie, and loss rates (%) of medical models vs. their corresponding base models across all (model pair, QA dataset) combinations. Win rate refers to the proportion of (model pair, QA dataset) combinations where a medical model shows a statistically significant improvement. We do observe that on textual medical knowledge QA tasks, medical LLMs outperform their base models in the SFT regime, although the improvements are marginal (1.6% average gain in accuracy; Section 5.1).
Table 1: Prior works overlook aspects critical to assessing the impact of medical DAPT on downstream zero-/few-shot performance, potentially leading to overly optimistic conclusions about its benefits. For each medical LLM/VLM included in our study (Table 2), we summarize the following details: (i) whether the medical model was compared against its base model; (ii) whether the sensitivity of the medical model and its base model to the choice of prompt format and few-shot examples was considered in their comparison; and (iii) whether statistical tests were performed to determine the significance of any measured performance improvement. ▲ indicates that the corresponding aspect was only partially considered (e.g., accounted for sensitivity to few-shot examples but not the prompt format).
Compared to Base Model? Accounted for Prompt Sensitivity? Tested for Statistical Significance?
Med42-v2 (Christophe et al., 2024b)
Med42-v1 (Christophe et al., 2024a)
OpenBioLLM (Pal and Sankarasubbu, 2024)
Clinical-Camel (Toma et al., 2023)
BioMistral (Labrak et al., 2024)
MediTron (Chen et al., 2023)
BioMedGPT-LM (Luo et al., 2023)
LLaVA-Med (Li et al., 2023a)
Med-Flamingo (Moor et al., 2023b)
Our Evaluation
@@
Overall, we find that both medical LLMs and VLMs show limited improvement over their base models, across all of the medical QA tasks and evaluation settings that we consider (Figure 2). In the zero-/few-shot prompting regime (Section 4), we find that all medical VLMs and the majority of medical LLMs fail to consistently outperform their base models across all datasets, including QA tasks focused on assessing medical knowledge and those based on real-world clinical notes. In the SFT regime (Section 5), we find that the medical LLMs overall do show statistically significant improvements on the medical knowledge QA tasks but not on the clinical note QA tasks, while the medical VLMs show little to no improvement on all of the visual medical QA tasks. Our findings suggest that a rigorous pairwise comparison of models, including tests for statistical significance, is essential to draw reliable conclusions about the performance benefits from medical DAPT, while these basic practices are not consistently adopted in the literature (Table 1).
@@
    We provide a comprehensive head-to-head comparison between state-of-the-art general-domain LLMs/VLMs and their medical DAPT counterparts on a wide range of medical QA tasks, to investigate the effectiveness of DAPT for medical specialization.
@@
    We find that in the SFT regime, all medical VLMs fail to show improvement, while medical LLMs show improvement on textual medical knowledge QA tasks but not on QA tasks based on real-world clinical notes (Section 5).
@@ 2 Related Work
More recent works suggest that decoder-based autoregressive LLMs and VLMs trained via medical DAPT also show strong performance on various downstream medical tasks. Medical LLMs such as MediTron (Chen et al., 2023), adapted from Llama-2 (Touvron et al., 2023b); and BioMistral (Labrak et al., 2024), adapted from Mistral-7B-Instruct-v0.1 (Jiang et al., 2023); perform well on knowledge-intensive QA tasks derived from medical licensing and academic exams (Jin et al., 2020; Pal et al., 2022; Hendrycks et al., 2021) and PubMed abstracts (Jin et al., 2019). Medical VLMs such as LLaVA-Med (Li et al., 2023a), adapted from LLaVA (Liu et al., 2023); and Med-Flamingo (Moor et al., 2023b), adapted from Open-Flamingo (Awadalla et al., 2023); also perform well on visual medical QA tasks based on radiology (Lau et al., 2018; Liu et al., 2021) and pathology images (He et al., 2020) and academic exams (Yue et al., 2024). These encouraging results have established DAPT as a go-to approach for training medically specialized LLMs and VLMs, a conclusion that we re-examine across different evaluation settings (zero-/few-shot prompting, SFT) and across both knowledge-intensive and clinically relevant QA tasks.
@@ 3 Experimental Setup
To investigate the effectiveness of medical DAPT in improving (i) zero-/few-shot and (ii) SFT performances, we compare ten medical LLMs and two medical VLMs against their general-domain counterparts in pairs on 22 textual medical QA datasets (8 based on clinical notes) and 8 visual medical QA datasets, respectively (Figure 2). The models in each pair are exactly identical in model architecture and scale, and their only difference lies in whether they were additionally pretrained on medical data. We also note that while some of the datasets considered contain both closed-ended questions (with objective ground-truth answers) and open-ended questions, we focus our evaluations on the former, where an objective, quantitative assessment of medical knowledge and reasoning capabilities is possible. To ensure the reproducibility of our results, we open-source the source code used for all of our evaluations described below via our GitHub repository333github.com/taekb/eval-medical-dapt.
LLM Llama-3-70B-Instruct Med42-v2-70B Medical QA Datasets (e.g., MedQA, MedMCQA)
Medical Instruction 120k (Altaf, 2023)
OpenGPT (OpenChat) (Wang et al., 2024)
StackExchange (Lambert et al., 2023)
Medical Flashcards (Han et al., 2023)
Llama-3-70B-Instruct OpenBioLLM-70B Undisclosed
Llama-2-70B MediTron-70B Clinical Practice Guidelines (e.g., CDC, WHO)
PubMed Articles (S2ORC; Lo et al., 2020)
Llama-2-70B Clinical-Camel-70B ShareGPT
20k PubMed Articles Published Before 2021
Random 4k Subset of MedQA (Jin et al., 2020)
Llama-2-70B Med42-v1-70B Medical QA Datasets (e.g., MedQA, MedMCQA)
OpenGPT (OpenChat) (Wang et al., 2024)
StackExchange (Lambert et al., 2023)
Medical Flashcards (Han et al., 2023)
CORD-19 (Wang et al., 2020)
Llama-3-8B-Instruct Med42-v2-8B Medical QA Datasets (e.g., MedQA, MedMCQA)
Medical Instruction 120k (Altaf, 2023)
OpenGPT (OpenChat) (Wang et al., 2024)
StackExchange (Lambert et al., 2023)
Medical Flashcards (Han et al., 2023)
Llama-3-8B OpenBioLLM-8B Undisclosed
Llama-2-7B MediTron-7B Clinical Practice Guidelines (e.g., CDC, WHO)
PubMed Articles (S2ORC; Lo et al., 2020)
Mistral-7B-Instruct-v0.1 BioMistral-7B PubMed Articles (PMC Open Access Subset)
Llama-2-7B-Chat BioMedGPT-LM-7B PubMed Articles (S2ORC; Lo et al., 2020)
VLM LLaVA-v0-7B LLaVA-Med-7B PubMed Articles (PMC-15M; Zhang et al., 2023)
Open-Flamingo-9B Med-Flamingo-9B Medical Textbooks (MTB; Moor et al., 2023b)
PubMed Articles (PMC-OA; Lin et al., 2023)
Models (Table 2).
For medical LLM evaluations, we consider the following model families: Med42-v2 (Christophe et al., 2024b), Med42-v1 (Christophe et al., 2024a), OpenBioLLM (Pal and Sankarasubbu, 2024), MediTron (Chen et al., 2023), Clinical-Camel (Toma et al., 2023), BioMistral (Labrak et al., 2024), and BioMedGPT-LM (Luo et al., 2023). For medical VLM evaluations, we consider LLaVA-Med (Li et al., 2023a) and Med-Flamingo (Moor et al., 2023b). For all models, we use the checkpoints made available via HuggingFace. In all zero-/few-shot prompting experiments, we generate predictions from each model via (i) greedy decoding (i.e., sampling with temperature $T=0$) and (ii) constrained decoding. For constrained decoding, we constrain the token vocabulary to be one of the answer choice letters (e.g., one of [“A”, “B”, “C”, “D”] for a four-choice QA dataset) and treat the answer choice with the highest token probability as a given model’s prediction. For the SFT experiments, we only generate predictions via greedy decoding, as it best reflects the setup used for fine-tuning (Section 3.2).
@@
For textual medical knowledge QA, we use MedQA (Jin et al., 2020), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019), and MMLU-Medical (Hendrycks et al., 2021) for evaluation. MMLU-Medical refers to a subset of MMLU corresponding to 9 subjects related to medicine: anatomy, clinical knowledge, college biology, college medicine, high school biology, medical genetics, nutrition, professional medicine, and virology. For MedQA, we use the official train–validation–test splits as provided through BigBio (Fries et al., 2022b). We note that MedQA has two versions, one with four answer choices per question and the other with five, and we use both for evaluation. For MedMCQA, which does not have a public test set, we follow the approach taken by Wu et al. (2024) and Labrak et al. (2024), taking a random 80–20 train–validation split of the official training set and using the official validation set for testing. For PubMedQA, we follow Singhal et al. (2023a), using the 211k artificially generated QA samples for training, and taking a 50–50 split on the 1k expert-labeled examples. For MMLU-Medical, we use the official split as provided. Meanwhile, as the Med42 models—Med42-v2-70B, Med42-v2-8B, and Med42-v1-70B—were trained on all of the QA datasets mentioned above as part of DAPT, we exclude these models from our discussion on textual medical knowledge QA.
@@
As strong performance on medical knowledge QA tasks may not necessarily imply high clinical utility (Wornow et al., 2023), we also perform evaluations with the following textual QA datasets based on clinical notes: MedNLI (Romanov and Shivade, 2018), EHRNoteQA (Kweon et al., 2024), the 2008 i2b2 Obesity Comorbidity Detection Challenge dataset (Uzuner, 2009) (4 classification tasks), the CASI Clinical Acronym Sense Disambiguation dataset (Moon et al., 2014), and the MIMIC-III Clinical Acronym Sense Disambiguation dataset (Johnson et al., 2016; Adams et al., 2020). For MedNLI, we frame each natural language inference task as a three-way closed-ended QA task, where the answer to each question is one of [“entailment”, “contradiction”, and “neutral”], and use the official train–validation–test splits provided via PhysioNet (Goldberger et al., 2000) under a data use agreement. For EHRNoteQA, we use the “Level 1” subset, which only contains QA pairs with clinical notes at most 3k tokens in length, to account for the limited context window sizes of the LLMs that we evaluate, and take a random 60–20–20 split of the data to obtain the train, validation, and test sets. For the 2008 i2b2 Obesity Comorbidity Detection Challenge dataset, we follow the preprocessing steps in Ceballos-Arroyo et al. (2024) to obtain 4 binary classification (i.e., two-way multiple-choice QA) datasets, each focused on detecting the presence of (i) asthma, (ii) coronary artery disease (CAD), (iii) diabetes, and (iv) obesity from a given clinical note. We then select the QA pairs with clinical notes at most 3k tokens in length, as done for EHRNoteQA. As the preprocessing pipeline by Ceballos-Arroyo et al. (2024) outputs a train–test split, we take a random 80–20 split on the training set to obtain the train and validation sets. For the CASI and MIMIC-III sense disambiguation datasets, we follow the preprocessing steps detailed in Adams et al. (2020) for consistency with prior works (Agrawal et al., 2022).
Visual Medical QA Datasets.